Grade 11General Physics

The internal and external diameters of a hollow cylinder are measured with the help of a Vernier Calipers.
Their values are 4.23 ± 0.01cm and 3.87 ± 0.01cm respectively. Then the thickness of the wall of the cylinder is

Sara
5 Years agoGrade 11

1 Answer

Vikas TU
5 Years ago
Dear student 
Thickness =4.23−3.87=0.18
Error =0.01+0.01=0.02
As error always gets added
So,
Thickness =0.18±0.02 cm.